A fraction becomes 1/2 if 5 is subtracted from its numerator and 3 is subtracted from its denominator however if we divide the numerator by 2 and add 7 to the denominator or the fraction becomes 1/4 find the fraction?
Answer by Boreal(15235)

x/y is fraction
(x-5)/(y-3)=1/2
cross multiply
2x-10=y-3
2x-7=y
------------------
(x/2)/(y+7)=1/4
cross multiply
2x=y+7
y=2x-7
-----------------
any x will work, although the fraction is undefined for x=3.5
try 1
y will be -5
(1-5)/-5-3)=-4/-8=1/2
(1/2)/2=(1/4)
-------------
Try x=-2, y=-11
-2-5/-11-3=-7/-14=1/2
(-2/2)/-11+7=-1/-4=1/4
The fraction is
x/2x-7
Try 3.5, which would make this fraction's denominator 0; y=0
-1.5/-3=1/2
1.75/7=1/4